OpenStudy (anonymous):

help please.. i know its not written right...write the square root in form of asqrtb where a and b are integers and b has the least value..28sqrt..

OpenStudy (anonymous):

write the square root in form of \[\sqrt[a]{b}\] where a and b are integers and b has the least value. the number is \[\sqrt{28}\]

OpenStudy (jdoe0001):

$$ \huge a\sqrt{b}\ \ or\ \ \sqrt[a]{b} $$

OpenStudy (anonymous):

it would be the first one. sorry.

OpenStudy (jdoe0001):

$$ 28 \implies 2*2*7 \text{, after simplifying} \implies 2^2*7\\ \sqrt{28} \implies \sqrt{2^2*7} $$ can you get it from there?

OpenStudy (anonymous):

let me try hold on

OpenStudy (anonymous):

\[2\sqrt{14}\]

OpenStudy (anonymous):

no thats not right hold on

OpenStudy (jdoe0001):

well, $$ \sqrt{28} \implies \sqrt{2^2*7} \implies \sqrt{2^2}*\sqrt{7} $$

OpenStudy (anonymous):

\[2\sqrt{7}\]

OpenStudy (jdoe0001):

:D

OpenStudy (anonymous):

im still wrong?

OpenStudy (jdoe0001):

nope, that's correct

OpenStudy (anonymous):

yay thanks!

OpenStudy (jdoe0001):

yw